Question 2 Tries remaining: 2 Points out of 7.70 Suppose the demand for jackets was given by: Op- 220 -P. The supply of jackets is given by: Qs 10P-770. Suppose the price was $195 per jacket. Calculate whether there is a surplus or shortage of jackets at that price and the quantity of jackets associated with the surplus or shortage. Enter your response as Qs - Negative numbers indicate shortage. Positive numbers indicate surplus. (Round to the nearest two decimal places if necessary.) Flag questionE Answer: Check
Answer #1

Surplus = Qs - Qd

= (10P - 770) - (220 - P)

= 10P - 770 - 220 + P

= 11P - 990

= 11(195) - 990

= 2145 - 990

= 1155